OpenStudy (anonymous):

To the nearest whole number, what is the surface area of a sphere with a volume of 140 cm3?

OpenStudy (anonymous):

V=4/3pir^3

OpenStudy (anonymous):

140=4/3pir^3 3/4*140=r^3

OpenStudy (anonymous):

{x=-2.358846990158267+4.085642834235041*i, x=-2.358846990158267-(4.085642834235041*i), x=4.717693980316533} is what my calculator got lol

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

Yeah lol, no idea

OpenStudy (anonymous):

alright so \[V=\frac{4}{3}\pi r^3\] if you know V = 140 \[140=\frac{4}{3}\pi r^3\] multiplying by the reciprocal \[.75*140=\pi r^3\] \[105=\pi r^3\] divide by pi and take the cube root to find an aproximate radius

OpenStudy (anonymous):

Thanks!